OpenStudy (anonymous):

A race car accelerates on a straight road from rest to speed of 180 km/hr in 25s . Assuming uniform acceleration of the car through out find the find the distance covered in this time .

OpenStudy (anonymous):

\[v=180 Km/hr=50 m/s\] \[v=aT \rightarrow a=\frac{ v }{ T }\] \[d=\frac{ 1 }{ 2 }aT^2=\frac{ 1 }{ 2 }\frac{ v }{ T } T^2=\frac{ vT }{2 }=\frac{ 25 \times 50 }{2 }=625m\]

OpenStudy (anonymous):

f=180*1000/60*60*25=2m/s^2 s=0+1/2*2*25^2=2500m=2.5km

OpenStudy (anonymous):

@marufkhan, I am afraid that: s=0+1/2*2*25^2=625m and not 2.5Km

OpenStudy (anonymous):

f=180*1000/60*60*25=2m/s^2 s=0+1/2*2*25^2=625m=.625km @CarlosGp@ tnx .. i did wrong calculation. sorry for that.